The series combination of 5 F and 20 F capacitors is connected in parallel to 4 F capacitor. What is the energy stored in the combination of capacitors?

1
Expert's answer
2022-03-21T12:42:27-0400

The net capacitance of system

"C_{net}=\frac{C_1C_2}{C_1+C_2}+C_3"

"C_{net}=\frac{5\times20}{5+20}+4=8\rm \: F"

The energy stored in the combination of capacitors

"W=\frac{C_{net}V^2}{2}"